OpenStudy (anonymous):

Factor completely. 4x^2+8x+4

OpenStudy (anonymous):

(2x+2)^2

OpenStudy (anonymous):

this can be put in the form (2x)^2 + 2(2x)(2) + (2)^2 which is of the form x^2 +2xy + y^2 = (x+y)^2 = (2x+2)^2 = (2x+2)(2x+2) = 4(x+1)(x+1)

OpenStudy (anonymous):

yep hark

OpenStudy (anonymous):

tks jimmy :))